OpenStudy (anonymous):

Write the equation in exponential form. Log4 1/16=-2 A. B. 42 = 16 C. D.

OpenStudy (jdoe0001):

hint: \(\bf log_{\color{red}{ a}}{\color{blue}{ b}}=y \iff {\color{red}{ a}}^y={\color{blue}{ b}}\qquad\qquad % exponential notation 2nd form {\color{red}{ a}}^y={\color{blue}{ b}}\iff log_{\color{red}{ a}}{\color{blue}{ b}}=y \\ \quad \\ thus \\ \quad \\ log_{\color{red}{ 4}}{\color{blue}{ \frac{1}{16}}}=-2\implies ?\)

OpenStudy (anonymous):

Umm can you walk me threw it i suck at this

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, is quite clear from there

OpenStudy (anonymous):

I dont get it tho

OpenStudy (jdoe0001):

look at the template above

OpenStudy (anonymous):

16 1/2 =4?

jhonyy9 (jhonyy9):

do you see there above what will be ,,a" and what will be ,,b" and what ,,y" ?

OpenStudy (anonymous):

I think so

jhonyy9 (jhonyy9):

no there are 4 1/16 -2 can you clearifie what will be a and b and y ?

jhonyy9 (jhonyy9):

so little help a=4 can you continue it ?

OpenStudy (anonymous):

B 1/16 Y -2 ???

jhonyy9 (jhonyy9):

yes - so now just use the formula what have wrote there above @jdoe0001

jhonyy9 (jhonyy9):

a^y = b

OpenStudy (anonymous):

4^-2=1/16

jhonyy9 (jhonyy9):

yes but do you see it there on these choices ?

OpenStudy (anonymous):

Yes thank you so much!!!!!

jhonyy9 (jhonyy9):

np ws my pleasure good luck bye